In a study of television viewing, the mean number of television program they watched during daytime was 7. A survey was conducted on the random sample of 25 households and found that the mean number of television program they watched during daytime was 5 with a standard deviation of 1.5. Test the hypothesis at 10% level of significance.

Parameter: _______________________________________

Statistic: ________________________________

Null Hypothesis: ____________________________________________________

Alternative hypothesis: _____________________________________________


1
Expert's answer
2022-05-08T14:12:55-0400

Parameter: "\mu," the mean number of television program they watched during daytime

Statistic: "t-" statistic, two-tailed test

Null Hypothesis: "H_0:\mu=7"

Alternative hypothesis: "H_a:\mu\not=7"

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=7"

"H_a:\mu\not=7"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


2. Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=24" degrees of freedom, and the critical value for a two-tailed test is "t_c = 1.710882."


3. The rejection region for this two-tailed test is "R = \{t: |t| > 1.710882\}."


4. The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{5-7}{1.5/\sqrt{25}}\approx-6.6667"


5. Since it is observed that "|t| =6.6667>1.710882=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=-6.6667" is "p=0.000001," and since "p= 0.000001<0.10=\alpha," it is concluded that the null hypothesis is rejected.


6.Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 7, at the "\alpha = 0.10" significance level.